40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 20 02:31:09 2015, ericq, Update, LSC, PRFPMI is back lock.pdf
    Reply  Tue Aug 25 04:15:51 2015, ericq, Update, LSC, PRFPMI is back 
       Reply  Tue Aug 25 16:09:54 2015, ericq, Update, IOO, IMC Tweak imcTweak.pdfCARM_TF.pdf
          Reply  Fri Aug 28 19:05:53 2015, rana, Update, IOO, IMC Tweak 
             Reply  Tue Sep 1 11:56:56 2015, ericq, Update, IOO, IMC loop shapes IMCshapes_Aug31_2015.pdfIMC_Aug31_2015.zip
                Reply  Wed Sep 2 01:31:22 2015, ericq, Update, IOO, IMC loop shapes imc_olg_trend.pdfimc_xover_trend.pdf
       Reply  Thu Aug 27 04:23:04 2015, ericq, Update, LSC, PRFPMI is back 
          Reply  Fri Aug 28 02:20:35 2015, Ignacio, Update, LSC, PRFPMI and MCL FF carmin.pngcarmout.png
Message ID: 11528     Entry time: Tue Aug 25 04:15:51 2015     In reply to: 11518     Reply to this: 11529   11534
Author: ericq 
Type: Update 
Category: LSC 
Subject: PRFPMI is back 

More PRFPMI locks tonight. Right now, it's been locked for 22+ minutes, though with the PRMI still on 3F signals. I think the MC2/AO crossover needs some reshaping; there's a whole bunch of noise injected into CARM around 600 Hz, which is where the two paths differ by 180deg. (Addendum: broke lock at ~27 minutes, 4:16AM)

For most of this lock, sensing matrix excitations have been running for daytime analysis. 

The nominal IMC loop gain / EOM crossover were making the AO path very marginal. I've adjusted the nominal settings and autolocker scripts. 

There was some weird behavior of X green PDH earlier... Broadband RIN seen in ALS-TRX, coherent with the DC output of the beat PD, so really on the light. I fiddled with the end setup, and it mostly went away, though I didn't intentionally change anything. Disconcerting. 

ELOG V3.1.3-